OpenStudy (anonymous):

If f(x) = 3a|4x – 4| – ax, where a is some constant, find f ′(1). 0 not enough information e 1

jimthompson5910 (jim_thompson5910):

hint: \[\Large f(x) = |x|\] \[\Large f'(x) = \frac{|x|}{x}\] this is a rule you memorize but you can derive this rule by breaking up |x| into +x and -x and deriving each piece (think of it as a piecewise function). notice how f ' (x) is undefined when x = 0, so that means f(x) is not differentiable at x = 0.

OpenStudy (anonymous):

so what should i get as the answer?

OpenStudy (anonymous):

0 right?

OpenStudy (anonymous):

@jim_thompson5910

jimthompson5910 (jim_thompson5910):

one sec

jimthompson5910 (jim_thompson5910):

do you see how I got f ' (x) ?

OpenStudy (anonymous):

yes @jim_thompson5910

OpenStudy (shadowlegendx):

Nice c;

OpenStudy (anonymous):

so what should i get as the answer?

jimthompson5910 (jim_thompson5910):

So if f(x) = 3a*|4x-4| - ax, then \[\Large f'(x) = \frac{3a|4x-4|}{4x-4}-a\]

jimthompson5910 (jim_thompson5910):

what is the value of f ' (1) ?

OpenStudy (zarkon):

that derivative needs a little work

OpenStudy (zarkon):

I'm referring to what @jim_thompson5910 wrote

jimthompson5910 (jim_thompson5910):

oh right, forgot the chain rule

OpenStudy (zarkon):

yes

jimthompson5910 (jim_thompson5910):

if f(x) = 3a*|4x-4| - ax, then \[\Large f'(x) = 4*\frac{3a|4x-4|}{4x-4}-a\] \[\Large f'(x) = \frac{12a|4x-4|}{4x-4}-a\]

OpenStudy (anonymous):

what does that give me as a result? Please hurry i'm at the library and I have to go in 5 minutes Im running out of time :(

OpenStudy (anonymous):

@Zarkon @jim_thompson5910

jimthompson5910 (jim_thompson5910):

you can work on it when you get more time you need to find f ' (1) based on what I posted this here: \[\Large f'(x) = \frac{12a|4x-4|}{4x-4}-a\]

OpenStudy (anonymous):

It's due very soon

OpenStudy (anonymous):

so I j ust plug in 1 for x?

OpenStudy (anonymous):

@jim_thompson5910

OpenStudy (anonymous):

that would give me 11a

OpenStudy (anonymous):

Or 0, right?

jimthompson5910 (jim_thompson5910):

into f ' (x) and not f(x)

jimthompson5910 (jim_thompson5910):

what is 4x - 4 when x = 1

OpenStudy (anonymous):

0

jimthompson5910 (jim_thompson5910):

can you divide by zero?

jimthompson5910 (jim_thompson5910):

can you say something like 2/0 ?

OpenStudy (anonymous):

no

OpenStudy (anonymous):

@jim_thompson5910

OpenStudy (anonymous):

wouldn't that just be equal to 0?

jimthompson5910 (jim_thompson5910):

2/0 is undefined

jimthompson5910 (jim_thompson5910):

so \[\Large f'(x) = \frac{12a|4x-4|}{4x-4}-a\] \[\Large f'(1) = \frac{12a|4*1-4|}{4*1-4}-a\] \[\Large f'(1) = \frac{12a|0|}{0}-a\] \[\Large f'(1) = \ \text{Undefined}\]

OpenStudy (anonymous):

that's not an answer option: my options are: a)0 b)e c)not enough information d)1

OpenStudy (anonymous):

@jim_thompson5910

OpenStudy (zarkon):

for \(a\ne0\) jim_thompson5910's answer is correct what happens when \(a=0\)

OpenStudy (anonymous):

the answer is = to 0?

OpenStudy (zarkon):

right...so the answer depends on \(a\)...and we don't know the value of \(a\)

OpenStudy (anonymous):

so the answer is not enough information?!

OpenStudy (zarkon):

right...we need to know if a is zero or not. So we don't have enough info

OpenStudy (anonymous):

:) ok THANK YOU!!!